OpenStudy (sissyedgar):

♥ I'm not too good at solving for the variable when there's a variable on both sides.♥ Enter the solution to the equation in the box. 5(x−8)=7(x−4) x = ?

OpenStudy (518nad):

hi okay first expand the brackets

OpenStudy (518nad):

5*x - 5*8 = 7*x-7*4

OpenStudy (bthompson361):

distribute first then isolate x

OpenStudy (bthompson361):

5x-40=7x-28

OpenStudy (bthompson361):

subtract 5x from each side -40=2x-28 add 28 to each side -12=2x divide by 2 on each side x=-6